OpenStudy (anonymous):

Given the ordered pair (3, y), what value of y would make the ordered pair a solution of the equation 4x − 2y = 24?

OpenStudy (ajprincess):

Put 3 in place of x in 4x-2y=24 and solve for y. Can u do it @DamnImBad?

OpenStudy (anonymous):

4(3)-2=10

OpenStudy (ajprincess):

4x − 2y = 24 4(3)-2y=24 Nw solve for y.

OpenStudy (ajprincess):

@DamnImBad

OpenStudy (anonymous):

-6

OpenStudy (ajprincess):

ya